# Factuality Beyond Coherence: Evaluating LLM Watermarking Methods for Medical Texts

### v1

@@ 5.4 Risk of Factuality Corruption (RQ4)
Entropy distribution by disease entity  As shown in Figure 4, the reference token entropy for disease entities [neumann-etal-2019-scispacy] generally aligns with the second entropy range (2–4) in Figure  3. Within this range, we observe noticeable shifts in median token entropy across disease entities shown in the plot, such as pain, infection, and cancer. These shifts suggest a degradation in the model’s contextual confidence when handling critical domain-specific terms. It also indicates that the more uncertain a model is on a disease entity, the more likely it is to hallucinate about that entity. This intuition will further be observed in a more detailed entity-level comparison.
@@
[…]
@@ 6 Conclusion
This study includes a human evaluation to validate the alignment between GPT-Judger assessments and human judgments, restricted to the Question Answering (QA) task. However, our objective was not to establish comprehensive human evaluation across all tasks, but to assess the reliability of our proposed evaluation workflow, which we demonstrate through a strong correlation with human judgments in the QA setting. Additionally, we identified contributing factors to factuality degradation like entropy distribution shifts and entity hallucinations as important observations, however other potential factors likely exist beyond our current analysis. The connections between these mechanisms and various watermarking approaches represents a promising area for further research toward developing more domain-sensitive watermarking techniques.
@@ A.1 Model Analysis
Table 8 presents a comparison of watermarking methods across medical language models for text QA tasks: Meditron 7B [^6] https://huggingface.co/epfl-llm/meditron-7b, MedLlama-3 8B [^7] https://huggingface.co/johnsnowlabs/JSL-MedLlama-3-8B-v2.0 and BioMistral 7B [^8] https://huggingface.co/BioMistral/BioMistral-7B. High detection metrics (TPR=1, AUROC=1) can be seen on smaller models (Meditron 7B and BioMistral 7B), with a decrease observed in the larger MedLlama-3 8B model. As expected, quality metrics improve with model size, with MedLlama-3 8B showing substantially higher SimCSE scores compared to smaller models, and better task performance in all metrics.
@@ A.2 Hyperparameter Effect on Detection
KGW $\gamma=0.5,\delta=2.0$
SWEET $\gamma=0.5,\delta=2.0$, $\text{entropy\_threshold}=0.9$
DiPmark $\gamma=0.5,\alpha=0.45$
EXP-edit $\text{pseudo\_length}=256$
@@
Additionally, we conducted a systematic investigation into how hyperparameter values affect detection performance. First, we explored the impact of varying $\delta\in\{0.5,1,2\}$ and $\gamma\in\{0.1,0.25,0.5\}$ in logit-based watermarking, KGW, and SWEET, with results presented in Figure 5. For these experiments, we utilized BioGPT [^10] https://huggingface.co/microsoft/biogpt [luo2022biogpt] model, which is computationally less demanding than the larger models used in our main analysis, enabling more efficient hyperparameter exploration. Our findings demonstrate that for both KGW and SWEET schemes, larger values of $\delta$ and $\gamma$ correspond to improved detection rates. However, as discussed in § 5.1, there exists a fundamental trade-off between detection efficacy and the quality of generated content.
@@ B.1 GPT-Judger
[…]
@@ B.3.1 Text Completion
For this task, we utilize the HealthQA dataset [zhu2019hierarchical], which consists of medical question-answer pairs. We focus solely on the answer portions, selecting 200 answers that exceed 230 words in length. From each answer, we extract the last 230 words, using the first 30 words as the input prompt and tasking the model with generating the subsequent 200 words. This experimental design aligns with prior watermarking evaluation frameworks [kirchenbauer2023watermark], allowing for consistent comparison across domains.
@@ B.3.2 Question Answering
We employ the same HealthQA dataset [zhu2019hierarchical] for our QA task, but utilize both question and answer components. To maintain consistency with the Text Completion task, we select 200 data points containing questions of exactly 10 words and answers of fewer than 250 words. The questions (ending with ?) serve as input prompts without additional instructions, while the corresponding answers constitute the ground truth for evaluation.
@@ B.3.3 Summarization
For the Summarization task, we employ the MeQSum dataset [abacha2019summarization], which provides pairs of detailed clinical questions and their concise summaries. We construct model prompts by prefixing the clinical question text with Write a short question that summarizes this question: and appending Summarized Question: after the question text. To introduce diversity, we limit our selection to inputs with a maximum of 60 words and summaries containing at least 10 words, creating more variation in input-output length ratios compared to the other tasks.
@@ B.4.1 KGW [kirchenbauer2023watermark]
B.4.2 SWEET [lee-etal-2024-wrote]
SWEET improve KGW [kirchenbauer2023watermark] by distinguishing watermark applicable tokens, meaning it embed and detect watermarks only within tokens with high entropy.
@@ B.4.2 SWEET [lee-etal-2024-wrote]
B.4.3 EXP-edit [kuditipudi2024robust]
This method apply the watermarking via Exponential Minimum Sampling [aaronson2022watermarking] with slight modification of detection using Levenshtein Alignment.

### v2

@@ 5.4 Risk of Factuality Corruption (RQ4)
Entropy distribution by disease entity  As shown in Figure 4, the reference token entropy for disease entities Neumann et al. (2019) generally aligns with the second entropy range (2–4) in Figure  3. Within this range, we observe noticeable shifts in median token entropy across disease entities shown in the plot, such as pain, infection, and cancer. These shifts suggest a degradation in the model’s contextual confidence when handling critical domain-specific terms. It also indicates that the more uncertain a model is on a disease entity, the more likely it is to hallucinate about that entity. This intuition will further be observed in a more detailed entity-level comparison.
@@
Entity hallucinations  Inspired by LLMs factuality hallucination described by Li et al. (2024), which categorizes erroneous entities as entity-error hallucination, we analyze factuality corruption in watermarked medical texts. Using a disease entity recognizer Neumann et al. (2019), we measure Introduced entities (new disease entities not found in reference text) and Hallucination rate (percentage of semantically divergent entities). We compute pairwise contextual BERT embeddings with a cosine similarity Zhang et al. (2020) threshold of 0.6, which allows us to distinguish acceptable synonyms from true semantic shifts that corrupt the text’s medical context.
@@ 6 Conclusion
This study includes a human evaluation to validate the alignment between GPT-Judger assessments and human judgments, restricted to the Question Answering (QA) task. Our objective was not to establish comprehensive human evaluation across all tasks, but to assess the reliability of our proposed workflow, which we demonstrate through a strong correlation with human judgments in the QA setting. We also identified contributing factors to factuality degradation, such as entropy distribution shifts and entity hallucinations, while other potential factors likely exist beyond our analysis. The connections between these mechanisms and various watermarking approaches represent a promising area for further research toward developing more domain-sensitive watermarking techniques.
@@ A.1 Model Analysis
Table 8 presents a comparison of watermarking methods across medical language models for text QA tasks: Meditron 7B [^4] https://hf.co/epfl-llm/meditron-7b, MedLlama-3 8B [^5] https://hf.co/johnsnowlabs/JSL-MedLlama-3-8B-v2.0 and BioMistral 7B [^6] https://hf.co/BioMistral/BioMistral-7B. High detection metrics (TPR=1, AUROC=1) can be seen on smaller models (Meditron 7B and BioMistral 7B), with a decrease observed in the larger MedLlama-3 8B model. As expected, quality metrics improve with model size, with MedLlama-3 8B showing substantially higher SimCSE scores compared to smaller models, and better task performance in all metrics.
@@ A.2 Hyperparameter Effect on Detection
KGW
$\gamma=0.5,\delta=2.0$
SWEET
$\gamma=0.5,\delta=2.0$,
$\text{entropy\_threshold}=0.9$
DiPmark
$\gamma=0.5,\alpha=0.45$
EXP-edit
$\text{pseudo\_length}=256$
@@
Additionally, we conducted a systematic investigation into how hyperparameter values affect detection performance. First, we explored the impact of varying $\delta\in\{0.5,1,2\}$ and $\gamma\in\{0.1,0.25,0.5\}$ in logit-based watermarking, KGW, and SWEET, with results presented in Figure 5. For these experiments, we utilized BioGPT [^8] https://hf.co/microsoft/biogpt Luo et al. (2022) model, which is computationally less demanding than the larger models used in our main analysis, enabling more efficient hyperparameter exploration. Our findings demonstrate that for both KGW and SWEET schemes, larger values of $\delta$ and $\gamma$ correspond to improved detection rates. However, as discussed in § 5.1, there exists a fundamental trade-off between detection efficacy and the quality of generated content.
@@ B.1 GPT-Judger
The prompt for GPT-Judger formerly referred to the template of Singh and Zou (2024), where the aspects are tailored for the factuality dimension designed in Table 1. We add detailed questions following each aspect to further trigger deeper reasoning. These questions per aspect are also used as in the Human Evaluation questionnaire detailed in Appendix B.2. The example of prompt and response of GPT-Judger for Summarization task is available in Table 14.
@@ B.3.1 Text Completion
For this task, we utilize the HealthQA dataset Zhu et al. (2019), which consists of medical question-answer pairs. We focus solely on the answer portions, selecting 200 answers that exceed 230 words in length. From each answer, we extract the last 230 words, using the first 30 words as the input prompt and tasking the model with generating the subsequent 200 words. This experimental design aligns with prior watermarking evaluation frameworks Kirchenbauer et al. (2023), allowing for consistent comparison across domains.
@@ B.3.2 Question Answering
We employ the same HealthQA dataset Zhu et al. (2019) for our QA task, but utilize both question and answer components. To maintain consistency with the Text Completion task, we select 200 data points containing questions of exactly 10 words and answers of fewer than 250 words. The questions (ending with ?) serve as input prompts without additional instructions, while the corresponding answers constitute the ground truth for evaluation.
@@ B.3.3 Summarization
For the Summarization task, we employ the MeQSum dataset Abacha and Demner-Fushman (2019), which provides pairs of detailed clinical questions and their concise summaries. We construct model prompts by prefixing the clinical question text with Write a short question that summarizes this question: and appending Summarized Question: after the question text. To introduce diversity, we limit our selection to inputs with a maximum of 60 words and summaries containing at least 10 words, creating more variation in input-output length ratios compared to the other tasks.
@@ B.4 Watermarking Methods
B.4.2 SWEET Lee et al. (2024)
SWEET improve KGW Kirchenbauer et al. (2023) by distinguishing watermark applicable tokens, meaning it embed and detect watermarks only within tokens with high entropy.
@@
B.4.3 EXP-edit Kuditipudi et al. (2024)
This method apply the watermarking via Exponential Minimum Sampling Aaronson and Kirchner (2022) with slight modification of detection using Levenshtein Alignment.